### v1

# Satellite Data Shows Resilience of Tigrayan Farmers in Crop Cultivation During Civil War

Abstract
The Tigray War was an armed conflict that took place primarily in the Tigray region of northern Ethiopia from November 3, 2020 to November 2, 2022. Given the importance of agriculture in Tigray to livelihoods and food security, determining the impact of the war on cultivated area is critical, but quantifying this impact was difficult due to restricted movement within and into the region due to conflict-driven insecurity and blockages. Using satellite imagery and statistical area estimation techniques, we assessed changes in crop cultivation area in Tigray before and during the war. Our findings show that cultivated area was largely stable between 2020-2021 despite the widespread impacts of the war. We estimated $1,132,000\pm 133,000$ hectares of cultivation in pre-war 2020 compared to $1,217,000\pm 132,000$ hectares in mid-war 2021. Comparing changes inside and outside of a 5 km buffer around conflict events, we found a slightly higher upper confidence limit of cropland loss within the buffer (0-3%) compared to outside the buffer (0-1%). Our results support other reports that despite widespread war-related disruptions, Tigrayan farmers were largely able to sustain cultivation. Our study demonstrates the capability of remote sensing combined with machine learning and statistical techniques to provide timely, transparent area estimates for monitoring food security in regions inaccessible due to conflict.
keywords:
Crop mapping, Conflict, Area estimation, Land use change, Food security, Smallholder farming, Machine learning
[…]
left=2.5cm,right=2.5cm,top=2.5cm,bottom=2.5cm
1 Introduction
Armed conflicts profoundly disrupt agriculture and food systems, jeopardizing food security and livelihoods for affected populations [1]. Direct impacts include damage to agricultural infrastructure, loss of labor and inputs, market dysfunctions, and prevention of farming activities [2, 3]. Indirect impacts include environmental degradation, uncontrolled resource exploitation, and displacement of farming communities [4]. Quantifying agricultural land use changes is critical for assessing the impact of war on food security in inaccessible conflict regions and targeting humanitarian relief efforts [5].
The Tigray War was an armed conflict primarily fought in the Tigray region of northern Ethiopia that lasted from November 3, 2020 to November 2, 2022. The conflict involved the Ethiopian federal government and Eritrea on one side and the Tigray People’s Liberation Front (TPLF) on the other. Figure 1 provides a map and timeline of conflict events recorded by the Armed Conflict Location and Event Data Project (ACLED) in 2020 and 2021 (the timeline considered for our study). Violence and destruction were widespread in the Tigray region, though most conflict events occurred in Southern, North Western, and Central Tigray. Most battles and other incidents occurred between November 2020 and March 2021, and reports of fighting decreased substantially after Ethiopia declared a unilateral cease-fire in June 2021 to allow farmers to till their land and humanitarian aid groups to enter.
(a) Map of conflict events in ACLED dataset in 2020 and 2021
(b) Timeline of events in ACLED database. The crop calendar and timeline of key political events are shown at the bottom, including the official start of the war in November 2020.
Figure 1: Map and timeline of conflict events from ACLED database
Agriculture is a critical industry in Tigray, providing employment for 80% of people and constituting 65% of the land area [6]. Disruptions to cultivation could have severe consequences for livelihoods and food security, thus quantifying the impact of the war on agriculture is crucial for targeting humanitarian relief to ensure the population’s basic needs are met. Many farmers reported delays in ploughing due to insecurity [3] and 353,000 people in Tigray were reportedly experiencing “catastrophic” levels of “acute food insecurity” in June 2021 [6]. Food insecurity in early 2021 was exacerbated by the loss of food stocks destroyed by armed forces during the start of the war in late 2020 and the blockage of food and humanitarian aid to the region [6]. Humanitarian aid was blocked or severely limited from the start of the conflict until April 2021, when aid organizations could deliver food aid for the first time since late 2020. In surveys conducted between April and May 2021, 50-70% of households in NorthWestern and Central Tigray reported that food assistance was their primary source of food [6]. The June 2021 cease-fire gave hope that farmers would be able to cultivate their fields in the 2021 main season, but there were grave concerns and uncertainty about food insecurity and the fate of food production in 2021.
Satellite remote sensing data can provide qualitative and quantitative insights about the impact of conflict on agricultural land use and production. This remote, real-time data source is especially critical for assessing impacts in regions where traditional ground-based surveys cannot be conducted due to insecurity, such as Tigray during 2021. Most studies using satellite data to assess war impacts on cultivation in other regions have found significant reductions in cultivated land due to war, ranging from approximately 15-30% across studies. Skakun et al. [7] found a 22% net loss of cropland in regions not occupied by the Ukraine government following the military conflict that began in 2014 in southeastern Ukraine, and that cropland losses were higher (46%) in areas in the buffer zone surrounding the conflict border. In a study of the more recent war in Ukraine that began in 2022, Ma et al. [8] detected concentrations of fallowed croplands in western Kherson and central Luhansk associated with conflict activities. Witmer [9] found that the average cultivated area decreased by 25% in the period after the Bosnian War between 1992-1995. In a study of the armed conflict in Mali, Boudinaud and Orenstein [10] found that 25% of areas in Mopti exhibited crop area losses in 2019 compared to pre-conflict years. Olsen et al. [11] concluded that cropland cultivation decreased by 16% from 2016 to 2018 as a result of armed conflict in South Sudan. Müller and Kummerle [12] and Peterson and Aunap [13] studied the impact of government and economic transitions in Romania after 1989 and Estonia between 1990-1993, respectively; while these transitions were not armed conflicts, they were tumultuous periods with disruptions to the agricultural sector that were associated with 32% of croplands being abandoned in Estonia and 21% abandonment in Romania. In the first (1994-1996) and second (1999-2009) Chechen wars, Yin et al. [14] found higher rates of cropland abandonment in areas closer to conflict events (45% abandonment within 1 km compared to 6% greater than 8 km from conflict events) as well as higher rates of re-cultivation in areas farther away from conflict events. While most studies have concluded net losses or negative impacts to agricultural production due to conflict, Li et al. [15] found that there was a net increase in rural areas of Syria during the civil war and Eklund et al. [16] reported a mixture of cropland expansion, abandonment, and other changes in areas of Syria and Iraq occupied by the Islamic State. NASA Harvest [17] found that despite fears of disruptions to wheat production in Ukraine driven by the Russia-Ukraine war, wheat production in Ukraine in 2023 was similar to pre-war production levels. In a study of cropland abandonment in 14 regions globally, Yin et al. [18] found that detecting cropland abandonment was particularly challenging in regions characterized by small field sizes and low-intensity farms such as Nepal and Uganda. Together, these studies highlight that the impacts of armed conflict on agricultural land use is highly variable in different contexts. Quantifying changes from satellite data remains challenging, particularly in smallholder farming contexts, and requires a mixture of evidence to shed light on inaccessible regions.
[…]
In this study, we used well-established cropland mapping and area estimation techniques leveraging remote sensing data, machine learning classifiers, and statistical estimators to estimate the change in cropped area in 2020 (pre-war) compared to 2021 (wartime) in Tigray. We used recommended practices for computing scientifically rigorous and transparent estimates of change in land area [20]. Our results provide evidence that complements previous studies of the impact of war on cultivation in Tigray [3, 19]. Our results show that cultivated area was largely stable between 2020 and 2021, with a relatively small amount of land attributed to cropland gain or loss. Comparing the change in cultivated area in a 5 km buffer of conflict events to the remainder of Tigray, we found a slightly higher upper limit of the confidence interval of the percentage of cropland loss within the conflict buffer (0-3%) compared to outside (0-1%). Our findings support the conclusion that Tigrayan farmers were largely able to cultivate crops at a scale comparable to pre-war years despite the widespread impact of the war on farming and other aspects of life.
2 Methods
2.1 Study area
[…]
2.2 Conflict data
We obtained a geographically and temporally referenced dataset of conflict events that occurred in Tigray in 2020 and 2021 from the Armed Conflict Location and Event Data Project (ACLED). ACLED compiles data from various sources including media reports, NGOs and international organizations, government reports and statements, social media and online platforms, academic literature, eyewitness accounts, and local sources. We included all event types provided in the ACLED dataset except “Peaceful Protests.” This resulted in a dataset of 831 events with multiple event types: 666 battles, 104 violence against civilians, 43 explosions/remote violence, 11 strategic developments, 5 riots, and 2 protests. Figure 1 shows a map and timeline of these events. From Figure 0(b), it can be seen that the majority of conflict events occurred prior to the start of the 2021 growing season. This plot also shows that Ethiopia’s declaration of an immediate, unilateral cease-fire occurred during the land preparation phase. However, some battles and other events were still reported after the cease-fire. Of the five admin2 zones in Tigray, the majority of the conflict events in the 2020-2021 time period were recorded in the Southern zone (247), followed by North Western (199), Central (196), Eastern (117), and Western (72).
(a) Process diagram for creating annual cropland maps
(b) Process diagram for estimating annual cropland area, change area, and map accuracy
Figure 2: Process diagrams for creating annual cropland maps and computing area estimates. Map images are for illustration purposes and do not represent actual results.
2.3 Cropland mapping
2.3.1 Classification model
Figure 1(a) summarizes the workflow we used to create annual cropland maps. We used the long short term memory (LSTM) recurrent neural network architecture from Kerner et al. [23] to create a binary cropland map for 2020 and 2021 separately. The LSTM architecture has one base hidden layer, which is followed by two classification heads: one two-layer head is used to classify training instances within Tigray, and one single-layer head is used to classify global training instances (outside Tigray). We used a batch size of 128 instances. We used early stopping with a patience of 10 epochs and proceeded with model weights from the training checkpoint with the highest F1 score evaluated on the validation set. Training concluded after 16 epochs for the 2020 model and 13 epochs for the 2021 model (one epoch represents one complete pass through the entire training dataset). We used the Adam optimizer with default parameters to tune weights during training with learning rate set to $0.001$. We set the $\alpha$ hyperparameter that balances the global and local samples in the loss function to $10$ [23]. We used dropout during training with a dropout probability of $0.2$. We used the trained model for each year to predict the full cropland map for Tigray in 2020 and 2021 using the OpenMapFlow python library [24]. Our code used to create the cropland maps in this study can be found at https://github.com/nasaharvest/crop-mask.
2.3.2 Cropland change mapping
We created a cropland change map for 2020-2021 by combining the cropland maps created for 2020 and 2021 separately. This resulted in a map with four change classes: stable non-crop, stable crop, crop gain, crop loss.
2.3.3 Satellite data
We used the pixel-level time series input data format from the CropHarvest database [25] for the LSTM classifier. The 12-month time series consists of aggregated observations in a given pixel over each month from February 1 to January 31 for the year of interest (2020 or 2021) from Sentinel-2, Sentinel-1, and the Shuttle Radar Topographic Mission (SRTM) Digital Elevation Model (DEM). Note that while CropHarvest includes meteorological data (precipitation and temperature) from the ERA5 dataset, we omitted this dataset because we found that it introduced artifacts in the predicted map. We used all bands from Sentinel-2 except B1 (coastal aerosol) and B10 (cirrus) and added NDVI (computed from B8 and B4). This resulted in 12 channels from Sentinel-2. We aggregated the 5-day observations to monthly by taking the least-cloudy pixel per 30-day window [26]. We used top-of-atmosphere observations (Level 1A) since this is required by the algorithm. We used the VV and VH channels from the Sentinel-1 synthetic aperture radar dataset which we aggregated to monthly by taking the median of observations over each month. We used the SRTM DEM to compute elevation and slope features. In total, our input data had 16 channels for each of 12 timesteps for a given pixel (note that elevation and slope are constant over all timesteps). All datasets were pre-processed, resampled to 10 m/pixel, and exported to Google Cloud Storage using Google Earth Engine. The preprocessing code for each dataset can be found at https://github.com/nasaharvest/openmapflow/tree/main/openmapflow/eo.
2.3.4 Labeling and dataset splits
Our training dataset consisted of approximately 100,000 total instances with binary labels of crop or non-crop (about 5,000 of which were in Tigray and thus used to train the local classification head). The training dataset was constructed primarily from publicly accessible data in the CropHarvest dataset (https://zenodo.org/doi/10.5281/zenodo.5021761). The validation and test datasets can be found at https://data.harvestportal.org/dataset/annual-and-change-maps-for-tigray-2020-2021. We created validation and test datasets for choosing hyperparameter settings and evaluating final performance metrics (respectively) using a random uniform sample of points in Tigray; we created independent samples for each year. Each point was labeled by the authors and other trained individuals using the Collect Earth Online platform using a combination of monthly PlanetScope (3-5 m/pixel) composites, monthly Sentinel-2 composites, Mapbox Satellite basemap, and other high-resolution images available from Google Earth. Annotators answered the question “Does this point contain active cropland?” for each point based on whether the point location visually resembled an agricultural field and showed signs of sowing, growing, and harvesting during the approximate crop calendar of Tigray. As in Skakun et al. [7], we defined cropland as “a piece of land that is sown/planted and harvestable at least once within the 12 months after the sowing/planting date.” Each point was labeled by at least two different annotators and we retained only the points for which both (or a majority of) labelers agreed. This resulted in a validation set of 440 labels (140 crop and 300 non-crop) for 2020 and 289 labels (93 crop and 196 non-crop) for 2021, and a test set of 425 labels (127 crop and 298 non-crop) for 2020 and 297 labels (109 crop and 188 non-crop) for 2021.
2.3.5 Post-processing
We obtained a small ground-truth survey dataset collected in Tigray during 2021 by University of Ghent and Mekelle University [27]. This dataset contains ground-truth labels of agricultural fields in Tigray with annotations of “cultivated” (127 fields) or “fallow” (34 fields). To better understand the difficulty of distinguishing cultivated from fallow fields (which may resemble cultivated fields due to the presence of weeds or other vegetation) in satellite observations, we plotted the 12-month time series profiles in each of the 12 Sentinel-2 channels used for our model input. Figure 3 shows the NDVI time series and 7 shows the time series for all Sentinel-2 bands. Figure 7 shows that the time series for all bands is visually very similar for the cultivated and fallow fields surveyed. This observation, in addition to visual inspection of predicted crop maps in Tigray, suggested that our predicted crop maps may overestimate cultivated area in Tigray by classifying some fallow fields as cultivated.
Figure 3: Left: 12-month time series of Sentinel-2 NDVI for ground-truth plots annotated as cultivated (“Crop”), fallow with weeds (“Fallow Weed”), or fallow with no vegetation at all (“Fallow None”). Solid lines depict the mean and shaded regions depict the standard deviation of the time series for all samples. Right: Visualizations of example ground-truth plots in PlanetScope September 2021 basemap. The plot labeled “Fallow with weeds or grass” has a similar appearance to fields labeled with crop types, while the field labeled “Fallow (no vegetaion at all)” is easier to distinguish. Note that our map classification used Sentinel-2 inputs (left-most time series) while our response design used PlanetScope data for photointerpretation of reference sample labels (similar to those shown on the right).
In Figure 3, some differentiation between cultivated and fallow fields can be seen in the NDVI time series: the peak of the NDVI time series appears substantially higher on average than the peak reached in fallow fields. This also makes intuitive sense: we would expect fallow fields to have lower biomass than cultivated fields, since the pattern of vegetation in fallow fields should be sparse and unmaintained. Thus, we hypothesized that for a given predicted cropland map, within the pixels predicted as cropland there will be a minority sub-population of pixels that reached lower peak NDVI than the majority of the pixels which may correspond to fallow fields (or other erroneous predictions). To identify and remove these potential false positives, we computed the mean and standard deviation of the peak NDVI value for all positive (cropland) predictions in a given map, and identified which of those pixels had NDVI above or below $n$ standard deviations from the mean (i.e., $\mu\pm n\sigma$). We then re-classified those “anomalous” positive pixels to the non-crop class. We evaluated the effect of this post-processing algorithm using $n\in[0,0.5,1.0,1.5,2.0,2.5,3.0,3.5]$ on the true positive rate (TPR) versus false positive rate (FPR) for the ground-truth dataset and validation dataset for Tigray in 2021. We found that a threshold of $n=3.5$ offered the optimal tradeoff between TPR and FPR for the ground-truth dataset, without significantly increasing the FPR for the validation dataset. Thus for our final cropland maps in 2020 and 2021, we applied this post-processing algorithm with $n=3.5$ in an effort to reduce false positive predictions.
2.4 Area estimation
Reliable area estimates cannot be obtained by pixel counting (i.e., summing the pixels in each class and multiplying by the area covered by each pixel) from maps that inevitably contain classification errors. Sample-based area estimates are considered good practice for producing unbiased estimates of land area [20]. Mapped areas obtained from pixel counting and sample-based areas may differ significantly [28]. We used the best practices outlined in Olofsson et al. [20] to compute unbiased area estimates using a reference sample stratified by the classification map classes. Figure 1(b) summarizes the workflow we used to estimate annual cropland area and change in cropland area between two years.
2.4.1 Sampling and response design
We created a stratified random sample of $10\times 10$-meter pixels using the four strata in the 2020-2021 cropland change map classes. We drew 799 samples total (sampling rate 0.00015% in Tigray) using the following allocation to the four strata: 131 for “stable planted,” 404 for “stable not-planted,” 134 for “planted gain,” and 130 for “planted loss.” We pre-allocated 100 samples to each change stratum and then allocated the remaining samples approximately proportionally to the map area for each strata to reduce the uncertainties in area estimates for these classes resulting from low sample sizes [20]. We labeled reference samples using visual interpretation of satellite imagery by trained individuals, using the same procedure and satellite data sources described in Section 2.3.4. Annotators answered two questions, “Does this point contain active cropland in 2020?” and “Does this point contain active cropland in 2021?”, for each point based on whether the point location visually resembled an agricultural field and showed signs of sowing, growing, and harvesting during the approximate crop calendar of Tigray. Each reference sample was independently labeled by at least two annotators. For samples that did not have unanimous label agreement for both questions, we convened expert annotators as a group to discuss and decide the label to assign for each year.
2.4.2 Cultivated area in Tigray and admin2 zones
We used the reference samples to compute confusion matrices between the map class and reference labels, which were then used to compute map accuracy, area estimates and uncertainties, and investigate changes between 2020-2021 that could be seen visually in satellite images [20]. We computed area estimates for the four change classes representing transitions from 2020 to 2021 as well as annual cropland area estimates for 2020 and 2021 separately. For the annual area estimates, we used the reference labels from each year provided by the two-year change reference samples. We computed the annual and change area estimates for the Tigray region as well as within each of the five admin2 zones in Tigray. For each zone, we clipped the classification map to the zone boundary and filtered the reference samples to use only the samples within a given zone to estimate area. The code used to compute annual area estimates is available at https://github.com/nasaharvest/crop-mask/blob/165d01cc3114427fac59f28ec2eeee5c97de25fc/notebooks/crop_area_estimation.ipynb. The code used to compute change area estimates is available at https://github.com/nasaharvest/crop-mask/blob/359708379fb7f064e1248ff35daaf56e27eee97a/notebooks/ethiopia_tigray_change_area_estimation.ipynb. The reference samples can be accessed via Github following the paths recorded in the change area estimation notebook.
2.4.3 Cultivated area in proximity to conflict events
To evaluate whether cropland change was concentrated in locations where conflict events directly occurred, we computed the four-class change area estimates inside and outside of a buffer surrounding locations of reported conflict events. Using the ACLED conflict event database described in Section 2.2, we used QGIS to create a 5 km buffer surrounding each event location and then dissolved the boundaries to create a single geometry. We chose 5 km as the buffer size based on previous studies that used similar buffer sizes to assess conflict impacts on agriculture [11, 7, 19]. We estimated cropland change inside this boundary by clipping the change map and filtering the reference samples to the buffer geometry. To estimate cropland change outside of the buffer zone, we created a new geometry representing the difference between the Tigray region boundary and the buffer zone boundary. We then estimated cropland change in the outside-buffer boundary by clipping the change map and filtering the reference samples to the geometry.
3 Results
3.1 Map classification and accuracy assessment
Figure 4 shows the annual cropland classification maps for 2020 and 2021 as well as the four-class change map of transitions between the 2020 and 2021 maps. Full resolution GeoTIFFs of maps are available at https://data.harvestportal.org/dataset/annual-and-change-maps-for-tigray-2020-2021.
(a) 2020 cropland map
(b) 2021 cropland map
(c) 2020-2021 change map
Figure 4: Annual cropland and change maps overlaid with admin2 zone boundaries (solid lines) and 5 km buffer around conflict events (dashed lines).
Table 1 and Table 2 report the overall accuracy, precision (user’s accuracy), recall (producer’s accuracy), and F1 score for the 2020 and 2021 maps, respectively. We computed these metrics using the independent test sets labeled for each year described in Section 2.3.4. We found F1 scores of $0.78\pm 0.11$ for the 2021 map and $0.69\pm 0.11$ for 2020. Overall accuracy was similar between the two years: $0.83\pm 0.02$ in 2020 and $0.84\pm 0.02$ in 2021. There was a moderate difference in precision (user’s accuracy) between 2020 ($0.75\pm 0.04$) and 2021 ($0.79\pm 0.04$) and larger differences in recall (producer’s accuracy) between 2020 ($0.65\pm 0.03$) and 2021 ($0.76\pm 0.03$). This is likely due to differences in climatic factors between the two years that result in different types of errors made by the classifier. Unbalanced precision scores between the two years may result in higher error rates when the maps are combined to form a change map, compared to maps with balanced precision scores.
Map Map Year F1 OA Recall (PA) Precision (UA) TN FP FN TP
WorldCover-v100 2020 $0.70\pm 0.11$ $0.84\pm 0.02$ $0.60\pm 0.03$ $0.84\pm 0.04$ 283 15 51 76
Digital Earth Africa 2019 $0.66\pm 0.12$ $0.81\pm 0.02$ $0.61\pm 0.03$ $0.72\pm 0.04$ 268 30 49 78
WorldCereal-v100 2020 $0.62\pm 0.11$ $0.82\pm 0.02$ $0.49\pm 0.03$ $0.84\pm 0.04$ 286 12 65 62
Copernicus 2019 $0.53\pm 0.12$ $0.72\pm 0.02$ $0.52\pm 0.04$ $0.54\pm 0.05$ 242 56 61 66
GLAD 2019 $0.51\pm 0.11$ $0.78\pm 0.02$ $0.38\pm 0.03$ $0.79\pm 0.05$ 285 13 79 48
Dynamic World 2020 $0.50\pm 0.11$ $0.78\pm 0.02$ $0.37\pm 0.03$ $0.76\pm 0.05$ 283 15 80 47
ASAP 2017 $0.45\pm 0.12$ $0.72\pm 0.02$ $0.39\pm 0.03$ $0.54\pm 0.05$ 255 43 77 50
Esri LULC 2020 $0.40\pm 0.10$ $0.76\pm 0.02$ $0.27\pm 0.02$ $0.77\pm 0.06$ 288 10 93 34
Harvest (ours) 2020 $0.69\pm 0.11$ $0.83\pm 0.02$ $0.65\pm 0.03$ $0.75\pm 0.04$ 270 28 45 82
Table 1: Performance metrics for 2020 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score).
Map Map Year F1 OA Recall (PA) Precision (UA) TN FP FN TP
WorldCover-v200 2021 $0.63\pm 0.12$ $0.76\pm 0.02$ $0.54\pm 0.03$ $0.75\pm 0.05$ 168 20 50 59
Copernicus 2019 $0.57\pm 0.09$ $0.67\pm 0.02$ $0.53\pm 0.02$ $0.61\pm 0.03$ 259 78 108 123
Digital Earth Africa 2019 $0.56\pm 0.13$ $0.70\pm 0.03$ $0.51\pm 0.04$ $0.61\pm 0.05$ 152 36 53 56
Dynamic World 2021 $0.53\pm 0.11$ $0.75\pm 0.03$ $0.39\pm 0.03$ $0.83\pm 0.05$ 179 9 66 43
WorldCereal-v100 2020 $0.53\pm 0.12$ $0.74\pm 0.03$ $0.40\pm 0.03$ $0.77\pm 0.06$ 175 13 65 44
GLAD 2019 $0.51\pm 0.12$ $0.72\pm 0.03$ $0.40\pm 0.03$ $0.71\pm 0.06$ 170 18 65 44
ASAP 2017 $0.49\pm 0.13$ $0.67\pm 0.03$ $0.42\pm 0.04$ $0.58\pm 0.06$ 154 34 63 46
Esri LULC 2021 $0.47\pm 0.12$ $0.72\pm 0.03$ $0.34\pm 0.03$ $0.77\pm 0.06$ 177 11 72 37
Harvest (ours) 2021 $0.78\pm 0.11$ $0.84\pm 0.02$ $0.76\pm 0.03$ $0.79\pm 0.04$ 166 22 26 83
Table 2: Performance metrics for 2021 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score).
Some prior studies used existing global land cover maps to derive a cropland map for change assessments instead of creating a custom classification map as in our study; for example, Weldegebriel et al. [19] used the WorldCover 10 m/pixel global land cover map for 2020 [29] to assess changes in cultivated area between 2020-2021 in Tigray. To assess the accuracy of our custom map classification compared to existing land cover maps, we evaluated multiple global land cover maps using our independent test set following the methodology described in Kerner et al. [30]. We only included maps that were created in 2017 or later to ensure temporal coherence: WorldCover [31, 32], Digital Earth Africa [33, 34], Copernicus [35, 36], Dynamic World [37], ASAP [38, 39], GLAD [40, 41], and Esri LULC [42, 43]. If multiple years were provided by a land cover product, we used the closest year to our reference data year. Of the compared maps, we found that WorldCover had the highest performance metrics overall; the F1 score for the WorldCover map in 2020 was comparable to our map but was lower for the 2021 map.
Year Southern Central Western Eastern North Western Tigray
2020 $227\pm 43$ $256\pm 48$ $327\pm 71$ $133\pm 37$ $282\pm 57$ $1,132\pm 133$
2021 $265\pm 45$ $241\pm 46$ $368\pm 76$ $136\pm 35$ $314\pm 60$ $1,217\pm 132$
Table 3: Annual area estimates for Tigray and each admin2 zone within Tigray (units are kha)
3.2 Annual area estimates
We report the annual cropland area estimates for 2020 and 2021 in Table 3. Figure 5 visualizes these estimates as a bar plot to illustrate the relative differences between each year. We found similar estimates for cultivated area in Tigray in 2020 and 2021, with a possible small net increase: $1,132\pm 133$ kha in 2020 and $1,217\pm 132$ kha in 2021 (1 kha = 1,000 ha). We found a similar trend for each admin2 zone, with the exception of the Central zone which showed a possible small net decrease. This net decrease in Central Tigray may be due to a higher concentration of conflict events recorded in the Central zone during the growing season [19]. Figure 6 shows the confusion matrices between the map and reference samples for Tigray. We provide the confusion matrices for admin2 zones in Figures 8-12. All confusion matrices are expressed in terms of proportion of mapped area based on good practice recommendations [20].
Figure 5: Annual area estimates for Tigray and each admin2 zone within Tigray (units are kha)
3.3 Change area estimates
Table 4 summarizes the area estimates for the four transition classes. We found that cropland gain accounted for between $70-160$ kha while cropland loss accounted for $19-79$ kha in Tigray. The stable transition classes accounted for the majority of the area in Tigray: $1,033-1,293$ kha of stable cropland and $3,798-4,054$ kha of stable non-cropland. Figure 5(a) shows the confusion matrix between the reference samples (stratified by the four change class strata) and the classified change map, expressed in terms of mapped area proportions. For Tigray, there was good agreement between the map and reference samples for the stable not planted class: of the area mapped as stable not planted, only a small fraction of the mapped area should have been mapped as stable planted according to the reference sample. There was also good agreement between the map and reference samples for the stable planted class, though approximately one third of the area mapped as stable planted should have been mapped as stable not planted according to the reference sample. This is likely due to the similarity in satellite time series between planted fields and other vegetated areas, which may include fallow fields in addition to non-agricultural areas. We used the same reference samples to compute class-wise accuracy metrics (overall accuracy, user accuracy/precision, producer accuracy/recall, true positive rate, and false positive rate) in Table 5.
The change estimates at the admin2 scale follow a similar trend as the admin1 (Tigray) scale. The majority of the area in each zone was estimated as stable non-crop followed by stable cropland. A relatively small amount of cropland gain and loss was estimated in each zone. However, the standard errors relative to the area estimates were much larger at the admin2 scale compared to those for Tigray due to smaller reference sample sizes. The standard error was highest for the change classes in Western Tigray. Figures 7(a)-11(a) provide the confusion matrices for each admin2 zone. Tables 7-11 report the accuracy metrics computed using the reference samples in each zone.
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 6: Confusion matrices between reference samples and Tigray change and annual cropland maps, expressed in terms of proportion of area as recommended in [20]
Class Southern Central Western Eastern North Western Tigray
Stable cropland $208\pm 43$ $226\pm 47$ $418\pm 80$ $108\pm 35$ $286\pm 58$ $1163\pm 130$
Stable non-crop $728\pm 47$ $775\pm 49$ $865\pm 78$ $473\pm 37$ $937\pm 59$ $3926\pm 128$
Cropland gain $53\pm 23$ $23\pm 17$ $28\pm 29$ $24\pm 14$ $21\pm 14$ $115\pm 45$
Cropland loss $25\pm 17$ $17\pm 13$ $10\pm 19$ $19\pm 16$ $12\pm 15$ $49\pm 30$
Table 4: Change area estimates for Tigray and each zone (admin2 level) within Tigray (units are kha)
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.69\pm 0.08$ $0.67\pm 0.06$ $0.53$ $0.07$
Stable non-crop $0.93\pm 0.03$ $0.81\pm 0.02$ $0.65$ $0.13$
Cropland gain $0.19\pm 0.07$ $0.55\pm 0.21$ $0.76$ $0.14$
Cropland loss $0.08\pm 0.05$ $0.59\pm 0.33$ $0.79$ $0.15$
Table 5: Accuracy metrics for four-class cropland change reference sample in Tigray. Overall accuracy is $0.77\pm 0.02$. See Tables 7-11 for admin2/zone-level metrics.
3.4 Cropland area changes in proximity to conflict events
Table 6 summarizes the change area estimates computed inside and outside of the 5 km conflict buffer. We reported the confidence interval in units of kha and as a percentage of the total area in each boundary to enable comparison between the two boundaries, since the total area outside the buffer is much larger than the area inside the buffer. We found that the percentage of total area estimated for each change class was similar within the confidence intervals of the estimates. However, our results show a higher upper boundary of cropland loss inside the conflict event buffer ($0-3\%$) compared to outside the buffer ($0-1\%$). This suggests that cropland loss may have been slightly more concentrated in the vicinity of conflict events compared to locations with greater distance from conflict events. Figure 13 shows the confusion matrix between the reference samples and the mapped areas inside and outside of the conflict event buffer. Tables 12 and 13 report the accuracy metrics.
Class Inside Buffer Outside Buffer
Stable cropland $318\pm 63$ ($19-29\%$) $838\pm 115$ ($18-24\%$)
Stable non-crop $943\pm 62$ ($67-76\%$) $2999\pm 113$ ($73-79\%$)
Cropland gain $37\pm 19$ ($1-4\%$) $77\pm 42$ ($1-3\%$)
Cropland loss $19\pm 17$ ($0-3\%$) $28\pm 23$ ($0-1\%$)
Table 6: Change area estimates (kha) inside and outside of a 5 km buffer around conflict events.
4 Discussion
4.1 Key findings and agreement with other studies
[…]
Our estimate of cropland loss in Tigray is much lower than cropland losses found in studies of some other conflict regions globally, which have been found to range from 15-30% [7, 9, 10]. Our findings further support the conclusion by Nyssen et al. [3] that Tigrayan smallholder farmers were largely able to sustain cultivation amidst violence. However, this may to some extent also reflect the difficulty of quantifying changes in small, heterogeneous farm landscapes like Tigray and confusion between planted and fallow fields, as discussed in Section 4.2 and Yin et al. [18].
4.2 Limitations
We followed best practices established by the scientific literature whenever possible, but there are some limitations inherent in our approach. One possible source of error in our estimates is in the reference labels used to estimate area. We tried to maximize the correctness and confidence of these reference labels as much as possible through our response design. We used high-resolution images from PlanetScope (3 m/pixel) and Google Earth ($<$1 m/pixel) when available. Each sample was analyzed and labeled independently by at least two different annotators; if there was any disagreement between annotators, we convened expert annotators to review and discuss all of the available evidence and come to a consensus label for the sample. Even with these precautions, it is possible that some samples were interpreted incorrectly due to the difficulty of differentiating between crop fields that are fallow and fields that are planted from satellite images, even in high-resolution (3 m/pixel) images (see Figure 3).
Another limitation of this study is that we used only two years for our change analysis: 2021 (mid-war) and 2020 (pre-war). Including additional pre-war years in the analysis would help establish a baseline of “nominal” area estimates to compare with mid-war 2021. It is also possible that in addition to the outbreak of war during the end of the harvest period in 2020, the COVID-19 pandemic may have had some impact on the 2020 growing season [44]. Nevertheless, we included only 2020 and 2021 in our study due to the significant resources required for analyzing two years and the precedence set by prior studies using similar statistical techniques that compared two years for change assessments (e.g., [20, 7, 11]).
Finally, the impacts of the Tigray war on crop cultivation in our study were limited to those that resulted in a binary change in cultivation status: either cultivated or not cultivated. This study did not account for other impacts such as population displacement, changes in crop types or cultivation practices, movement and access limitations, market dysfunction and loss of income, or destruction of food stocks and farming equipment that were reported in farmer surveys [6, 3]. While some studies have analyzed ploughing status and harvest status from satellite observations [3, 19], most of the other factors listed are difficult or impossible to determine from satellite observations, especially in the absence of extensive ground-truth data for calibration.
5 Conclusions
Despite widespread disruptions to agriculture, our remote sensing-based crop maps and area estimates suggest Tigrayan smallholder farmers were largely able to carry out crop cultivation activities in 2021 amidst the Tigray war. Our approach using satellite remote sensing, machine learning, and statistical techniques enables timely, accurate insights about agriculture in conflict regions where ground access is not possible using satellite data and statistical methods. Further studies could build on our methodology to incorporate additional years before and during the Tigray war to continue monitoring Tigray’s agriculture and food security situation. We made all code and data used in this study publicly accessible to enable transparency, reproducibility, and future work.
Acknowledgments
This work was supported by NASA under Award #80NSSC21K1511 and the NASA Harvest Consortium on Food Security and Agriculture (Award #80NSSC18M0039). We thank the Famine Early Warning Systems Network (FEWS NET) Science and Early Warning Teams for assistance with point labeling and discussions on early versions of results. We thank Liya Weldegebriel, Emnet Negash, and Jan Nyssen for sharing their concurrent findings, local expertise, and ground-truth data which helped strengthen this study.
Appendix A Sentinel-2 time series of cultivated and fallow fields
Figure 7: 12-month time series in each Sentinel-2 input channel for ground-truth plots annotated as cultivated (“Crop”) and fallow with weeds (“Fallow Weed”) or fallow with no weeds/bare soil (“Fallow None”). Solid lines depict the mean and shaded regions depict the standard deviation of the time series for all samples in each band.
Appendix B Admin2-scale confusion matrices
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 8: Confusion matrices for reference samples and change/annual cropland maps in Southern Tigray, expressed in terms of proportion of area as recommended in [20]
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 9: Confusion matrices for reference samples and change/annual cropland maps in Central Tigray, expressed in terms of proportion of area as recommended in [20]
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 10: Confusion matrices for reference samples and change/annual cropland maps in Western Tigray, expressed in terms of proportion of area as recommended in [20]
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 11: Confusion matrices for reference samples and change/annual cropland maps in Eastern Tigray, expressed in terms of proportion of area as recommended in [20]
(a) 2020-2021 change map
(b) 2020 map
(c) 2021 map
Figure 12: Confusion matrices for reference samples and change/annual cropland maps in North Western Tigray, expressed in terms of proportion of area as recommended in [20]
Appendix C Admin2-scale change map accuracy
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.42\pm 0.14$ $0.44\pm 0.10$ $0.32$ $0.11$
Stable non-crop $0.83\pm 0.05$ $0.75\pm 0.04$ $0.72$ $0.38$
Cropland gain $0.18\pm 0.12$ $0.25\pm 0.16$ $0.33$ $0.10$
Cropland loss $0.10\pm 0.09$ $0.27\pm 0.25$ $0.44$ $0.11$
Table 7: Accuracy metrics for four-class cropland change reference sample in Southern zone of Tigray. Overall accuracy is $0.65\pm 0.04$.
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.56\pm 0.19$ $0.32\pm 0.09$ $0.22$ $0.05$
Stable non-crop $0.82\pm 0.05$ $0.87\pm 0.03$ $0.81$ $0.46$
Cropland gain $0.13\pm 0.11$ $0.30\pm 0.27$ $0.56$ $0.11$
Cropland loss $0.22\pm 0.17$ $0.56\pm 0.39$ $0.71$ $0.06$
Table 8: Accuracy metrics for four-class cropland change reference sample in Central zone of Tigray. Overall accuracy is $0.73\pm 0.05$.
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.64\pm 0.13$ $0.83\pm 0.08$ $0.69$ $0.12$
Stable non-crop $0.91\pm 0.07$ $0.59\pm 0.05$ $0.41$ $0.11$
Cropland gain $0.08\pm 0.08$ $0.29\pm 0.36$ $0.60$ $0.18$
Cropland loss $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.00$ $0.23$
Table 9: Accuracy metrics for four-class cropland change reference sample in Western zone of Tigray. Overall accuracy is $0.66\pm 0.06$.
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.43\pm 0.21$ $0.28\pm 0.13$ $0.30$ $0.09$
Stable non-crop $0.85\pm 0.07$ $0.87\pm 0.03$ $0.79$ $0.34$
Cropland gain $0.36\pm 0.21$ $0.57\pm 0.31$ $0.73$ $0.09$
Cropland loss $0.07\pm 0.14$ $0.11\pm 0.21$ $0.17$ $0.08$
Table 10: Accuracy metrics for four-class cropland change reference sample in Eastern zone of Tigray. Overall accuracy is $0.74\pm 0.06$.
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.61\pm 0.15$ $0.39\pm 0.09$ $0.39$ $0.08$
Stable non-crop $0.84\pm 0.06$ $0.82\pm 0.03$ $0.69$ $0.34$
Cropland gain $0.19\pm 0.15$ $0.61\pm 0.34$ $0.62$ $0.08$
Cropland loss $0.02\pm 0.04$ $0.18\pm 0.35$ $0.33$ $0.16$
Table 11: Accuracy metrics for four-class cropland change reference sample in North Western zone of Tigray. Overall accuracy is $0.71\pm 0.05$.
Appendix D Conflict buffer confusion matrices
(a) Inside conflict buffer
(b) Outside conflict buffer
Figure 13: Confusion matrices for reference samples and change maps inside and outside 5km conflict event buffer, expressed in terms of proportion of area as recommended in [20]
Appendix E Conflict buffer change map accuracy
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.74\pm 0.13$ $0.66\pm 0.11$ $0.55$ $0.07$
Stable non-crop $0.92\pm 0.06$ $0.83\pm 0.04$ $0.63$ $0.10$
Cropland gain $0.30\pm 0.14$ $0.75\pm 0.30$ $0.87$ $0.15$
Cropland loss $0.11\pm 0.10$ $0.52\pm 0.45$ $0.67$ $0.15$
Table 12: Accuracy metrics for four-class cropland change reference sample inside 5km buffer of conflict events in Tigray. Overall accuracy is $0.78\pm 0.05$.
Metric Precision (UA) Recall (PA) True Positive Rate False Positive Rate
Stable cropland $0.66\pm 0.10$ $0.68\pm 0.07$ $0.51$ $0.06$
Stable non-crop $0.93\pm 0.03$ $0.80\pm 0.02$ $0.65$ $0.15$
Cropland gain $0.13\pm 0.07$ $0.43\pm 0.25$ $0.67$ $0.14$
Cropland loss $0.07\pm 0.05$ $0.66\pm 0.47$ $0.88$ $0.15$
Table 13: Accuracy metrics for four-class cropland change reference sample outside 5km buffer of conflict events in Tigray. Overall accuracy is $0.77\pm 0.03$.
qWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">